OpenStudy (anonymous):

b-9+6b=30 plz help me

OpenStudy (anonymous):

Simplifying b + -9 + -6b = 30 Reorder the terms: -9 + b + -6b = 30 Combine like terms: b + -6b = -5b -9 + -5b = 30 Solving -9 + -5b = 30 Solving for variable 'b'. Move all terms containing b to the left, all other terms to the right. Add '9' to each side of the equation. -9 + 9 + -5b = 30 + 9 Combine like terms: -9 + 9 = 0 0 + -5b = 30 + 9 -5b = 30 + 9 Combine like terms: 30 + 9 = 39 -5b = 39 Divide each side by '-5'. b = -7.8 Simplifying b = -7.8

OpenStudy (midhun.madhu1987):

 can u solve it from here..? combine the like terms..